### v1

Abstract
In multi-label emotion classification, particularly for low-resource languages like Arabic, the challenges of class imbalance and label correlation hinder model performance, especially in accurately predicting minority emotions. To address these issues, this study proposes a novel approach that combines stacked embeddings, meta-learning, and a hybrid loss function to enhance multi-label emotion classification for the Arabic language. The study extracts contextual embeddings from three fine-tuned language models—ArabicBERT, MarBERT, and AraBERT—which are then stacked to form enriched embeddings. A meta-learner is trained on these stacked embeddings, and the resulting concatenated representations are provided as input to a Bi-LSTM model, followed by a fully connected neural network for multi-label classification. To further improve performance, a hybrid loss function is introduced, incorporating class weighting, label correlation matrix, and contrastive learning, effectively addressing class imbalances and improving the handling of label correlations. Extensive experiments validate the proposed model’s performance across key metrics such as Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. The class-wise performance analysis demonstrates the hybrid loss function’s ability to significantly reduce disparities between majority and minority classes, resulting in a more balanced emotion classification. An ablation study highlights the contribution of each component, showing the superiority of the model compared to baseline approaches and other loss functions. This study not only advances multi-label emotion classification for Arabic but also presents a generalizable framework that can be adapted to other languages and domains, providing a significant step forward in addressing the challenges of low-resource emotion classification tasks.
keywords:
Multi-Label Emotion Classification, Class Imbalance, stacked embeddings, label correlation matrix and contrastive learning
1 Introduction
Social media platforms, particularly microblogging sites like Twitter, have revolutionized the way public sentiment and emotions are analyzed. Prior to social media, extracting valuable insights such as perceptions, opinions, and emotions from the public was a daunting task [1]. With millions of tweets posted daily, these platforms offer a rich source of data that reflects a wide range of human emotions and opinions [2, 3]. Understanding these emotional contents is crucial for providing context-related and personalized information, enhancing social movements, and improving various applications such as smart governance, education, and healthcare [3, 4].
Emotion Classification (EC) leverages Natural Language Processing (NLP) techniques to extract subjective information from textual data [5, 6]. This process is integral to multiple domains, including smart governance, business, and organizations, where it is applied to various data formats such as audio, text, and video. The theoretical foundation for EC is often grounded in psychological theories of emotion, such as Plutchik’s ”Wheel of Emotions” [7] and Ekman’s theory of basic emotions [8].
Automated EC, which includes Single-Label Emotion Classification (SLEC) and Multi-Label Emotion Classification (MLEC), has become essential due to the vast amount of text available online. MLEC is particularly advantageous as it identifies multiple emotions in a given text, addressing the limitations of SLEC which can only detect a single emotion per text.
In recent years, significant advancements in word embedding (WE) have contributed to the progress of MLEC. WE techniques represent words in dense, low-dimensional vectors that capture syntactic and semantic relationships, enabling better understanding and processing of human language by machines [9, 10]. However, Arabic text presents unique challenges for WE due to its complex morphology and diverse dialects [11].
Addressing the issues of dataset imbalance and class discrimination in MLEC is critical. Contrastive Learning (CL) has shown promise in learning effective representations by maximizing the similarity between related data samples and distinguishing dissimilar ones [12]. CL has achieved cutting-edge performance in capturing semantic representations of sentences [13] and addressing challenges in distinguishing discrete and valence dimension-based emotions [14].
Despite the progress in EC using various machine learning and deep learning approaches, Arabic MLEC faces challenges due to the small and imbalanced datasets available. Traditional loss functions like Binary Cross Entropy (BCE) are inadequate for handling class imbalance effectively [15]. To mitigate this, Focal Loss (FL) was introduced to focus on hard-to-classify samples [16]. Recent studies have proposed Multi-label Contrastive Focal Loss (MCFL), which combines FL with CL to enhance performance in imbalanced datasets [17, 18].
Deep learning models such as Bidirectional Long Short-Term Memory (Bi-LSTM) and Bidirectional Gated Recurrent Unit (Bi-GRU) have proven effective for NLP tasks. However, ensemble techniques, particularly stacking ensembles, offer improved predictions and generalization by combining outputs from multiple models [19]. Stacking ensembles utilize diverse models to enhance accuracy and robustness, making them suitable for complex tasks like Arabic MLEC.
This research proposes an MLEC model by employing stacked ensemble word embeddings derived from pre-trained language models including MARBERT, AraBERT, and ArabicBERT. Our objective is to harness the power of these models to generate high-contextual embeddings tailored for Arabic texts. Furthermore, we address the inherent challenges of dataset imbalance and class discrimination through the utilization of a hybrid loss function. This Arabic MLEC model aims to advance the field of Arabic MLEC by providing a robust and reliable model capable of enhancing emotion detection in Arabic tweets.
The main contributions of this study are outlined as follows:
  1. 1.
    Presented stacking approach for contextualized word embeddings by fine-tuning Arabic language models (MARBERT, AraBERT, and ArabicBERT) to facilitate high contextual embedding generation for Arabic MLEC.
  2. 2.
    Introduced a hybrid loss function by combining class weights, label correlation and contrastive learning, tailored for multi-label text classification, addressing challenges related to dataset imbalance and class discrimination.
  3. 3.
    Trained a second stage classifier as meta-learner which perform sequence learning using Bi-LSTM followed by multi-label classification learning using fully-connected network.
2 Related Work
Emotion classification (EC) is a pivotal area of research in natural language processing (NLP) with the goal of distinguishing between different emotional or psychological states expressed in text [20]. While substantial progress has been made in EC for high-resource languages like English, research focused on Arabic remains limited.
2.1 Multi-Label Emotion Classification for Arabic Language
Arabic MLEC has emerged as a significant research area attracting attention from various scholars aiming to address the unique challenges posed by Arabic text data. This section provides an overview of recent advancements in Arabic MLEC methodologies, with a focus on studies leveraging the SemEval-2018 task 1-Ec-Ar dataset [21] as a benchmark. Several studies have contributed to the advancement of Arabic MLEC, each employing diverse methodologies to tackle the complexities inherent in Arabic text emotion analysis. Notably, the SemEval-2018 dataset has served as a common benchmark, facilitating comparative analyses and informing the development of novel approaches.
Lin et al. [16] introduced contrastive learning techniques to enhance MLEC performance, achieving competitive F-macro and F-micro scores. Mansy et al. [22] explored the effectiveness of MARBERT, BiLSTM, and Bi-GRU models, demonstrating notable accuracy and precision in emotion detection tasks. Khalil et al. [23] leveraged Bi-LSTM AraVec/CBOW models, showcasing competitive performance in terms of Fmicro, precision, recall, and Jaccard accuracy.
Elfaik et al. [24] proposed a hybrid approach combining AraBERT word embeddings with attention-based LSTM and BiLSTM models, achieving high accuracy in MLEC tasks. Abdelali et al. [25] introduced the QARiB model and a hybrid feature model, yielding respectable F-macro scores in emotion classification. Additionally, Menai et al. [26] presented a hybrid feature-based model and deep feature-based model, demonstrating strong performance in terms of F-micro and Jaccard accuracy. Samy et al. [27] employed a context-aware GRU model, achieving competitive scores in emotion classification tasks.
The diverse range of methodologies and performance metrics employed in these studies [16, 22, 23, 24, 25, 26, 27] underscore the complexity of Arabic MLEC and the need for innovative approaches to address it. Furthermore, the utilization of the SemEval-2018 dataset as a benchmark has facilitated comparative analyses and enabled researchers to build upon existing methodologies.
2.1.1 Arabic Emotion Datasets
In addition to methodological advancements, the development of Arabic emotion datasets has played a crucial role in facilitating research in this domain. Table 1 provides a summary of existing Arabic emotion datasets, highlighting key attributes such as size, annotation type, accessibility, and emotion coverage. Despite the availability of these datasets, challenges related to accessibility and balance in emotion representation persist, underscoring the need for further efforts to address these limitations.
Table 1: Summary of Arabic Emotions Datasets
Authors/Year Dataset Size Type/Task Emotions
Althobaiti, (2023) ArPanEmo 11,128 DA/SLE Disgust, Happiness, Sadness, Confusion, Surprise, Anticipation, Optimism, Pessimism, Neutral
Shakil et al. (2021) AEELex 35,383 DA/SLE Anger, anticipation, disgust, fear, joy, sadness, surprise, trust
Azad et al. (2021) ExaACE 20,050 DA/MLE Anger, Anticipation, Love, Fear, Disgust, Joy, Sadness, Acceptance, Surprise, Neutral
Teahan (2019) IAEDS 1,366 DA/SLE Anger, Disgust, Happiness, Fear, Surprise, Sadness
Mohammad et al. (2018) SemEval 2018 4,381 DA/MLE Anger, Anticipation, Disgust, Fear, Joy, Love, Optimism, Pessimism, Sadness, Surprise, Trust, Neutral
ElBeltagy et al. (2018) ETDA 10,065 DA/SLE Joy, Sadness, Anger, Love, Fear, Sympathy, Surprise, None
Mohammad et al. (2018) SemEval 2018 5,600 DA/SLE Joy, Anger, Sadness, Fear
Alhuzali et al. (2018) LAMA-DIST 7,268 DA/SLE Anger, Anticipation, disgust, Fear, Joy, Sadness, Surprise, Trust
Abdul-Mageed et al. (2016) DINA 3,000 DA/SLE Anger, Disgust, Fear, Happiness, Sadness, Surprise
[…]
Therefore, the advancements in Arabic MLEC methodologies and the availability of benchmark datasets have significantly contributed to the progress of research in this field. However, continued efforts are required to overcome existing challenges and further enhance the effectiveness of Arabic MLEC systems.
2.2 Contextual Embeddings with Pre-Trained Language Models
Language processing is a complex task, requiring computational techniques to convert words into numerical formats that machines can understand. This process, known as Word Embedding (WE) or Word Representation, is critical in NLP activities, allowing texts to be represented in vector spaces [28]. WEs play a vital role in various NLP tasks, including machine translation, part-of-speech tagging, sequence labeling, named entity recognition, speech processing, and text classification [29, 30]. These embeddings capture the semantic meaning of words within text sequences, assigning numerical representations to words with similar meanings.
The effectiveness of supervised technologies in EC depends significantly on the features used in classification tasks [29]. Traditional WE techniques include frequency-based methods like Bag of Words (BoW), N-grams, and Term Frequency–Inverse Document Frequency (TF-IDF), as well as neural network-based methods like word2vec [31], Global Vectors for Word Representation (GloVe) [32], and FastText [33]. These earlier models, however, struggled to capture complex semantic dependencies and contextual information, often representing words with single, context-independent vectors [34, 35].
Recent advancements in NLP have led to the development of context-aware Pretrained Language Models (PLMs) such as Bidirectional Encoder Representations from Transformers (BERT) [36]. These models, utilizing transfer learning, have significantly improved performance across various NLP tasks. Unlike earlier WE models, PLMs like BERT excel in capturing complex semantic dependencies and contextual information by using transformer architecture for sub-word learning [36].
Numerous studies have compared the efficiency of different WE techniques across various domains. Al-Twairesh et al. [35] highlighted that while static models have shown progress, they suffer from contextual information deficiencies and out-of-vocabulary (OOV) issues. David et al. [37] compared static WE techniques using RNN and CNN for text classification, finding that FastText outperformed GloVe. Moreover, reducing batch size improved accuracy, demonstrating the significance of training parameters in model performance.
The stacking of WE techniques involves combining multiple embeddings to leverage their strengths. According to Muromägi et al. [38], stacked WE techniques can cancel out random noise from individual models and reinforce useful patterns. This approach is particularly beneficial in languages with limited textual data. Stacked WE techniques have shown improved performance over single models across various NLP tasks [39].
In the context of Arabic NLP, several BERT models such as AraBERT [40], ArabicBERT-Large [41], and MARBERT [42] have been developed. These models, trained on large-scale Arabic corpora, have demonstrated impressive performance in tasks like sentiment analysis and text classification [43]. AraBERT has shown superior performance in Arabic language understanding, while MARBERT, which focuses on dialectal Arabic, captures diverse Arabic dialects effectively [44].
Systematic reviews and comparative studies have underscored the effectiveness of BERT models in Arabic text classification. Alammary et al. [45] reviewed 66 studies, highlighting the efficiency of these models in various NLP tasks. Berrimi et al. [46] compared existing Arabic text classification approaches, supporting the use of BERT models for improved performance [47].
Despite the notable performance of individual BERT models, the stacking of these models has shown further improvements. Studies such as those by Kamr et al. [48] and Chouikhi et al. [49] have demonstrated that ensemble BERT models outperform single BERT techniques. However, no research has yet applied stacking techniques combining AraBERT, MARBERT, and ArabicBERT for feature extraction in Arabic Multilingual Emotion Classification (MLEC), despite the promising potential indicated by individual model performances.
The literature indicates that contextual embeddings using PLMs like BERT have revolutionized NLP by providing more accurate and context-aware word representations. The stacking of WE techniques further enhances model performance, particularly in languages with limited textual data.
2.3 Loss Functions for Imbalanced Data
Dataset imbalance, particularly prevalent in MLEC, presents significant challenges in achieving balanced and accurate model performance. According to Padma et al. [50] and Kumar et al. [51], a dataset is considered balanced when it contains an equal number of data points for all classes; otherwise, it is deemed imbalanced. This imbalance is especially noticeable in publicly available datasets. To address this, researchers [52, 53, 54] have proposed various loss functions designed to mitigate the effects of class imbalance, such as the Multi-label Focal Loss which addresses class imbalance by focusing more on hard-to-classify, minority class examples and reducing the emphasis on well-classified, majority class examples. In multi-label settings, it applies this weighting independently to each label, improving performance across both majority and minority classes.
Binary Cross Entropy (BCE) loss is a commonly used loss function for multi-label text classification tasks, including MLEC. However, BCE is highly susceptible to label imbalance, often influenced by the prevalence of dominant classes or negative instances [52]. To address this, Focal Loss (FL) was proposed by Lin et al. [55] to reshape the standard cross-entropy loss, placing more emphasis on hard and misclassified samples. Dong [53] highlights FL’s efficacy in focusing on difficult instances more effectively than classic cross-entropy loss. Jiang et al. [54] assert that while FL was initially applied in computer vision, it also shows reputable performance in text mining. Consequently, due to their success in other domains, loss functions like FL have been adapted for use in NLP.
Researchers such as Li et al. [34] and Sun et al. [56] have applied these adapted loss functions to NLP tasks. For instance, Sun et al. [56] applied FL to an information extraction model analyzing biomedical literature, reporting effective results in addressing imbalance. Similarly, Zhu et al. [57] utilized FL in speech emotion recognition, improving performance on the Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset.
While FL has demonstrated success in various tasks, its effectiveness is more pronounced in single-label classification. According to Ridnik et al. [58], treating positive and negative samples equally in multi-label classification, as FL does, can accumulate greater loss gradients from negative samples, reducing the importance of rare positive examples. This limitation has led to the development of Multi-label Focal Loss (MFL) to better accommodate multi-label classification tasks, including MLEC.
Deng et al. [59] applied MFL to address class imbalance in MLEC on the RenCECps and NLPCC2018 datasets, achieving better performance than state-of-the-art methods. Similarly, Bendjoudi et al. [60] used MFL in their multi-label emotion recognition model, achieving remarkable results and outperforming state-of-the-art methods for minority classes.
Contrastive learning (CL) has also been explored as a method for handling data imbalance. CL learns generalized representations by capturing similarities within a class and contrasting them with other classes [61]. This approach addresses the fine-grained feature spaces of similar emotions, enhancing generalization ability [62]. Researchers have proposed various CL methods to improve emotion discrimination, and these methods have shown promise in handling imbalanced data.
Combining Focal Loss and Contrastive Learning has led to the development of Multi-label Contrastive Focal Loss (MCFL). This combination leverages the strengths of both methods: FL’s focus on hard instances and CL’s ability to differentiate between similar and dissimilar data points. Studies by Chen et al. [18] and Zheng et al. [17] in computer vision have shown that MCFL outperforms state-of-the-art methods in imbalanced datasets. Lin et al. [16] further explored the efficiency of CL in multi-label text classification, including Arabic MLEC, using BERT-base-Arabic for fine-tuning. Their study proposed five novel multi-label contrastive losses, with the Stepwise Label Contrastive Loss (SLCL) achieving the highest FMacro and Strict Contrastive Loss (SCL) achieving the highest FMicro and Jaccard Score. These studies collectively highlight the effectiveness of combining loss functions to minimize the effects of class imbalance, allowing the model to efficiently learn multi-label classification. This research gap presents an opportunity to formulate a joint loss function that combines contrastive learning with variants addressing class imbalance, thus justifying the need for a hybrid loss function.
3 Methodology
This study adopts a systematic paradigm which views computer science programs as entities analogous to mental processes, integrating both formal deduction and scientific experimentation to acquire a priori and posteriori knowledge [63]. This approach is particularly prevalent in the field of artificial intelligence, ensuring rigorous and systematic investigation. The research focuses on the implementation of Arabic MLEC. The framework designed to achieve the research objectives is depicted in Figure 1. This framework outlines the procedural steps and methodological tools used, including data pre-processing, model development, and evaluation metrics. Guided by the scientific paradigm, the approach systematically address the challenges associated with Arabic MLEC, validating the findings through both theoretical and empirical means. This alignment with established practices in AI research ensures the robustness and credibility of the study’s results.
Figure 1: Framework of the proposed approach depicting various stages
3.1 Dataset Description
Utilizing existing datasets for text emotion classification is common among researchers. However, many online datasets suffer from significant class imbalances, impeding the performance of machine learning (ML) and deep learning (DL) models [64]. Such imbalances often lead models to favor majority classes, resulting in suboptimal predictions for minority classes.
In response to this issue, this study aims to address class imbalance in existing datasets, with a particular focus on the SemEval-2018 Task 1-Ec-Ar dataset curated by Mohammad et al. [21]. This dataset has been extensively employed in prior research on MLEC and provides Arabic text samples with annotated emotion labels, making it a valuable resource for developing and testing advanced MLEC models.
The dataset comprises eleven emotion classes, including anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust [21]. The dataset was collected using the Twitter API, with search queries aligned to tweets containing emotional words. Annotation reliability was determined using Best Worst Scaling (BWS), with each tweet labeled with ’1’ indicating the presence of the corresponding emotion class and ’0’ indicating its absence. Table 2 provides some sample instances containing Tweets and the multi-label emotion class identified for each Tweet.
Table 2: Tweets along with their multi-label emotion classes
Despite its usefulness, the SemEval-2018 Task 1-Ec-Ar dataset suffers from imbalanced emotion label distributions, a common challenge in real-world datasets. Figure 2 provides distribution of instances in each emotion class indicating class imbalance. This imbalance presents a significant obstacle for ML and DL models, particularly in accurately predicting underrepresented emotion classes.
Figure 2: Distribution of instances in each emotion category
To mitigate this challenge, this study proposes the utilization of an optimized loss function specifically designed to address class imbalance issue. By focusing on this aspect, the research aims to enhance the performance of MLEC models on the SemEval-2018 Task 1-Ec-Ar dataset, thereby facilitating more balanced and accurate emotion classification. This approach contributes to the development of more robust ML and DL models and enhances the reliability and applicability of emotion classification in Arabic text.
3.2 Pre-Processing
The preprocessing of text data is a crucial step in developing robust machine learning and deep learning models for NLP tasks such as MLEC. The preprocessing involved several steps to handle various text formatting issues, particularly focusing on the treatment of punctuation, emojis, emoticons, and normalization of Arabic text. Figure 3 provides a flowchart depicting the steps and their sequence to perform preprocessing of the text.
First, both English and Arabic punctuation marks were replaced with whitespace. This ensures that punctuation does not interfere with the tokenization process. Replacing punctuation helps in maintaining the integrity of the words and prevents any misinterpretation during the model training phase.
Figure 3: Framework of the proposed approach depicting various stages
Next, emojis and common emoticons were replaced with their corresponding textual tokens based on commonly available emoji and emoticon dictionaries. This approach preserves the emotional context conveyed by these symbols, which are prevalent in informal text. By converting emojis and emoticons to textual tokens, the preprocessing retains valuable emotional cues that might otherwise be lost if left untreated or removed from the data. This step ensures that the sentiment and emotional nuances embedded in the original text are captured and utilized effectively during model training.
The text was then stripped of English letters and numbers, focusing solely on the Arabic content. This step is essential for eliminating any extraneous characters that could skew the analysis towards non-Arabic language patterns. Similarly, Arabic numbers were also replaced with whitespace to ensure that the numerical data does not affect the linguistic model.
Normalization of Arabic text was another critical step. This process involved standardizing different forms of the same letter and removing diacritics. Normalization ensures that variations in spelling or character forms do not lead to misclassification or redundancy in the dataset. It helps in treating different forms of the same word as a single entity, thereby enhancing the consistency of the text data.
Specific characters such as dashes and decorative symbols were replaced with whitespace. These characters can often appear in social media texts and might not carry significant meaning related to the content. Removing them helps in focusing on the actual textual information. Similarly, backslashes were replaced with whitespace to prevent any potential issues in text parsing. Repeated characters were reduced to a single occurrence. This is particularly important in languages like Arabic, where character repetition can occur frequently for emphasis. Reducing repeated characters helps in normalizing the text and reducing redundancy. Extra whitespaces were removed to ensure that the text data is clean and compact. Additionally, standalone single characters were eliminated as they often do not contribute significant meaning and can be considered noise in the text data.
Further, specific patterns such as words starting with [Uncaptioned image] were addressed. The replacement of such patterns ensures that common prefixes are treated uniformly. This helps in maintaining the grammatical and contextual integrity of the sentences.
Finally, after these preprocessing steps, the text data was ready for extraction of word embeddings using pre-trained models from Hugging Face, specifically AraBERT, ArabicBERT and MARBERT. These models leverage state-of-the-art language representations and are well-suited for processing Arabic text, ensuring that the input data for the machine learning and deep learning models is of high quality and consistency. This comprehensive preprocessing pipeline is crucial for training effective models for Arabic MLEC, enhancing the reliability and accuracy of emotion detection in Arabic text.
3.3 Stacked Embeddings Extraction
To generate highly contextualized embeddings from Arabic text for multi-label emotion classification (MLEC), a stacked embedding extraction approach is introduced. This method leverages three pre-trained language models specifically tailored for the Arabic language: ArabicBERT, MarBERT, and AraBERT. Figure 4 illustrates the process of extracting embeddings from the preprocessed text.
3.3.1 Fine-Tuning the Language Models
The fine-tuning process is performed using the pre-processed SemEval-2018 dataset, which contains Arabic tweets annotated with multiple emotions. During preprocessing, the text is cleaned, tokenized, and formatted to match the input requirements of the pre-trained models. Special tokens like [CLS] and [SEP] are added to mark the beginning and end of each sequence.
Once prepared, the dataset is used to fine-tune ArabicBERT, MarBERT, and AraBERT. Fine-tuning involves adjusting the pre-trained weights to the multi-label emotion classification task. The models’ final classification layers are replaced with task-specific dense layers suitable for handling multiple labels. This modification allows the models to learn precise representations of emotion-laden Arabic text.
3.3.2 Embedding Extraction and Stacking
After fine-tuning, contextualized embeddings are extracted from the hidden layers of the models. These embeddings, either from the final layer or an aggregation of layers, now contain task-specific information relevant to the emotions in the dataset. By capturing both syntactic and semantic properties, these embeddings encapsulate rich contextual information specific to the Arabic language and the classification task.
Instead of using individual embeddings from each model, a stacked embedding approach is employed. Embeddings from ArabicBERT, MarBERT, and AraBERT are concatenated to create a comprehensive representation of the input text. This concatenation leverages the unique strengths of each model, providing a richer and more nuanced understanding of the text. The resulting stacked embeddings serve as highly contextualized inputs, which are then passed to a subsequent meta-learner.
The process of extracting embeddings and their subsequent concatenation is outlined in Algorithm 1. In this algorithm, embeddings are first extracted from the fine-tuned models and then concatenated to form the stacked embeddings.
Algorithm 1 Stacked Embeddings for Emotion Classification
0:  $models$: Pre-trained models (ArabicBERT, MarBERT, AraBERT)
0:  $T$: Preprocessed Arabic text (shape: [num_samples, sequence_length])
0:  $Y$: Emotion labels (shape: [num_samples, num_emotions])
0:  $E_{\text{stacked}}$: Stacked embeddings (shape: [num_samples, embedding_dim $\times$ 3])
1:  $T_{\text{processed}}\leftarrow\text{preprocess}(T)$
2:  for $model\in[\text{ArabicBERT},\text{MarBERT},\text{AraBERT}]$ do
3:     $model\leftarrow\text{fine\_tune}(model,T_{\text{processed}},Y)$
4:  end for
5:  $E_{\text{ArabicBERT}}\leftarrow\text{extract\_embeddings}(\text{ArabicBERT},T_{\text{processed}})$
6:  $E_{\text{MarBERT}}\leftarrow\text{extract\_embeddings}(\text{MarBERT},T_{\text{processed}})$
7:  $E_{\text{AraBERT}}\leftarrow\text{extract\_embeddings}(\text{AraBERT},T_{\text{processed}})$
8:  $E_{\text{stacked}}\leftarrow\text{concatenate}([E_{\text{ArabicBERT}},E_{\text{MarBERT}},E_{\text{AraBERT}}])$
9:  $Classification_{\text{Output}}\leftarrow\text{emotion\_classifier}(E_{\text{stacked}})$
10:  return  $E_{\text{stacked}}$
3.4 Meta-Learner
The meta-learner, designed as a robust recurrent neural network, is trained on the concatenated embeddings derived from the base models. This meta-learning component is tasked with integrating predictions or embeddings from the base models to refine overall performance.
Figure 4: Contextualized embeddings generation for stacking ensemble
The meta-learner architecture incorporates a Bidirectional Long Short-Term Memory (Bi-LSTM) network, which is renowned for its proficiency in capturing dependencies from both past and future contexts within textual data. The Bi-LSTM layer is meticulously configured with 25 units and includes dropout and recurrent dropout mechanisms set to 0.3. This strategic setup has the dual benefit of reducing overfitting tendencies while preserving critical information and fostering robust generalization capabilities.
Subsequent to the Bi-LSTM layer, a dense layer comprising 50 units with a Rectified Linear Unit (ReLU) activation function is employed for further processing. This facilitates feature extraction and refinement before the final classification layer. The output layer of the model comprises 12 units with a sigmoid activation function, accommodating the multi-label nature of the emotion classification task. Utilizing a sigmoid function, the model generates probabilities for each of the 12 emotion classes, signifying the presence or absence of each emotion. The architecture of the meta-learner is provided in Figure 5. Training of the model is executed using a proposed hybrid loss function, prioritizing minority classes and challenging samples to enhance model performance. The training parameters encompass a maximum sequence length of 32 tokens, a batch size of 32, and a learning rate of 0.001. The model undergoes training for 100 epochs to ensure convergence while simultaneously monitoring for overfitting and underfitting.
Figure 5: Architecture of the meta-learner used for classification
3.5 Loss Function
The performance of a deep learning model, especially in the context of multi-label classification tasks, is significantly influenced by the choice of the loss function. For the task of Arabic MLEC, various loss functions have been explored to handle class imbalance and leverage label correlations effectively. This section details the three loss function variations considered: class weights calculation, label correlation matrix, and contrastive learning.
3.5.1 Class Weights Calculation
Class imbalance is a prevalent issue in multi-label classification, where some classes may have significantly more instances than others as depicted in section 3.1. To address this, class weights are introduced into the loss function to give more importance to minority classes.
Let $w_{c}$ denote the weight for class $c$. The weight for each class can be calculated using the inverse frequency of the class as depicted in eq. (1).
$w_{c}=\frac{N}{C_{c}}$ (1)
  • where $N$ is the total number of instances, and $|C_{c}|$ is the number of instances in class $c$.
The weighted binary cross-entropy loss for each instance is then defined as eq. (2).
$L=-\sum_{c=1}^{C}w_{c}\left[y_{c}\log(p_{c})+(1-y_{c})\log(1-p_{c})\right]$ (2)
where:
  • $C$ is the number of classes.
  • $y_{c}$ is the ground truth label for class $c$ ($1$ if the instance belongs to class $c$, 00 otherwise).
  • $p_{c}$ is the predicted probability for class $c$.
This approach helps mitigate the impact of class imbalance by penalizing misclassifications in minority classes more heavily.
3.5.2 Label Correlation Matrix
In multi-label classification, the presence of certain labels can be correlated with others. A label correlation matrix can be utilized to incorporate these dependencies into the loss function.
Let $Y$ be the label matrix where each element $y_{ij}$ indicates the presence of label $j$ in instance $i$. The label correlation matrix $M$ is computed as eq. (3).
$M=\frac{Y^{T}Y}{N}$ (3)
The loss function is then adjusted to account for these correlations. In order to achieve this, the cross-entropy loss is modified to include a term that penalizes inconsistent predictions with respect to the label correlations as depicted in eq. (4).
$L=-\sum_{c=1}^{C}\left[y_{c}\log(p_{c})+(1-y_{c})\log(1-p_{c})\right]+\lambda\sum_{i=1}^{C}\sum_{j=1}^{C}M_{ij}(p_{i}-p_{j})^{2}$ (4)
where:
  • $\lambda$ is a regularization parameter controlling the strength of the correlation penalty.
  • $M_{ij}$ is the element of the label correlation matrix representing the correlation between labels $i$ and $j$.
This formulation encourages the model to make predictions that are consistent with the observed label correlations.
3.5.3 Contrastive Learning
Contrastive learning is a technique used to learn representations by comparing similar and dissimilar pairs of instances. In the context of multi-label classification, this approach can be adapted to improve the model’s ability to distinguish between different labels. The contrastive loss function is defined as eq. (5).
$L=\sum_{i=1}^{N}\sum_{j=1}^{N}\left[y_{ij}\cdot D(h_{i},h_{j})+(1-y_{ij})\cdot\max(0,m-D(h_{i},h_{j}))\right]$ (5)
where:
  • $y_{ij}$ is 1 if instances $i$ and $j$ share at least one label, and 0 otherwise.
  • $D(h_{i},h_{j})$ is the distance between the embeddings $h_{i}$ and $h_{j}$ of instances $i$ and $j$.
  • $m$ is a margin parameter that defines the minimum distance required between dissimilar instances.
By minimizing the loss described in eq. (5), the model learns to pull together the embeddings of instances with similar labels and push apart those with different labels, enhancing the discriminative power of the learned representations.
These three loss function variations—class weights calculation, label correlation matrix, and contrastive learning—provide robust mechanisms to address the challenges of multi-label emotion classification in imbalanced datasets. Class weights calculation ensures that minority classes receive higher importance during training, the label correlation matrix captures dependencies between labels to improve the prediction of rare classes, and contrastive learning enhances the model’s ability to differentiate between majority and minority classes, leading to improved overall performance.
3.5.4 Proposed Hybrid loss
The proposed hybrid loss function for multi-label classification integrates three key components: CL, LCM, and CW. These components are combined to create a unified loss function that addresses class imbalance and label correlations. By leveraging CL, LCM, and CW, the model is optimized to focus on hard-to-classify instances, capture label relationships, and balance class contributions, ensuring robust performance across all classes. Below is a brief explanation of each component:
  1. 1.
    Contrastive Learning (CL): Contrastive Learning helps the model focus on hard-to-classify instances by maximizing the distance between instances of different labels and minimizing the distance between similar ones. This mechanism enhances the model’s ability to differentiate between classes, improving classification accuracy, especially for minority classes.
  2. 2.
    Label Correlation Matrix (LCM): LCM captures the dependencies between labels, allowing the model to learn from label co-occurrence patterns. By accounting for correlations, LCM improves the model’s ability to predict labels that often appear together, making it more effective in handling complex label relationships in multi-label emotion classification.
  3. 3.
    Class Weights (CW): Class Weights mitigate the problem of class imbalance by assigning higher weights to minority classes and lower weights to majority classes. This ensures that the model pays more attention to underrepresented classes during training, leading to more balanced classification performance across all labels.
The pseudo-code 2 provides a detailed description of how these components are calculated and integrated into the hybrid loss function, ensuring that the model can efficiently generalize even in the presence of significant class imbalance and complex label correlations.
Algorithm 2 Hybrid Loss Function
0:  $y_{\text{true}}$: True labels (one-hot encoded, shape: [batch_size, num_classes])
0:  $y_{\text{pred}}$: Predicted logits (shape: [batch_size, num_classes])
0:  $embeddings$: Embeddings from the model (shape: [batch_size, embedding_dim])
0:  $class\_weights$: Weights for each class (shape: [num_classes])
0:  $label\_correlation\_matrix$: Correlation matrix for labels (shape: [num_classes, num_classes])
0:  $margin$: Margin parameter for contrastive learning
0:  $\alpha$, $\beta$, $\gamma$: Weighting coefficients for the loss components
0:  $hybrid\_loss$: Scalar value of the Hybrid loss
1:  $cl\_loss\leftarrow 0$
2:  for $i=1$ to $batch\_size$ do
3:     for $j=i+1$ to $batch\_size$ do
4:        if $y_{\text{true}}[i]==y_{\text{true}}[j]$ then
5:           $cl\_loss\leftarrow cl\_loss+\max(0,margin-\|embeddings[i]-embeddings[j]\|^{2})$
6:        else
7:           $cl\_loss\leftarrow cl\_loss+\|embeddings[i]-embeddings[j]\|^{2}$
8:        end if
9:     end for
10:  end for
11:  $cl\_loss\leftarrow cl\_loss/\left(\frac{batch\_size\times(batch\_size-1)}{2}\right)$
12:  $lcm\_loss\leftarrow 0$
13:  for $i=1$ to $num\_classes$ do
14:     for $j=1$ to $num\_classes$ do
15:        $lcm\_loss\leftarrow lcm\_loss+label\_correlation\_matrix[i][j]\times(y_{\text{pred}}[:,i]-y_{\text{true}}[:,i])\times(y_{\text{pred}}[:,j]-y_{\text{true}}[:,j])$
16:     end for
17:  end for
18:  $lcm\_loss\leftarrow lcm\_loss/num\_classes$
19:  $cw\_loss\leftarrow 0$
20:  for $i=1$ to $num\_classes$ do
21:     $cw\_loss\leftarrow cw\_loss+class\_weights[i]\times\text{CrossEntropyLoss}(y_{\text{true}}[:,i],y_{\text{pred}}[:,i])$
22:  end for
23:  $cw\_loss\leftarrow cw\_loss/num\_classes$
24:  $hybrid\_loss\leftarrow\alpha\times cl\_loss+\beta\times lcm\_loss+\gamma\times cw\_loss$
25:  return  $hybrid\_loss$
26:
27:
28:  function CrossEntropyLoss($y_{\text{true}}$, $y_{\text{pred}}$)
29:  return  $-y_{\text{true}}\times\log(y_{\text{pred}})-(1-y_{\text{true}})\times\log(1-y_{\text{pred}})$
3.6 Performance Evaluation Measures
To assess the performance of our MLEC model, we used several evaluation metrics. These metrics provide a comprehensive view of the model’s accuracy, ability to identify relevant instances, and its overall effectiveness in handling multi-label classification tasks.
3.6.1 Precision
Micro-averaged Precision measures the overall accuracy of the positive predictions across all labels. It is defined as the ratio of the sum of true positive predictions to the total number of positive predictions. The formula for micro-averaged precision is given by eq. (6):
$\text{Precision}=\frac{\sum_{i}TP_{i}}{\sum_{i}(TP_{i}+FP_{i})}$ (6)
where $TP_{i}$ is the number of true positives for the $i_{th}$ label and $FP_{i}$ is the number of false positives for the $i_{th}$ label.
3.6.2 Recall
Micro-averaged Recall measures the model’s overall ability to identify all relevant instances across all labels. It is defined as the ratio of the sum of true positive predictions to the total number of actual positives. The formula for micro-averaged recall is given by eq. (7)
$\text{Recall}=\frac{\sum_{i}TP_{i}}{\sum_{i}(TP_{i}+FN_{i})}$ (7)
where $TP_{i}$ is the number of true positives for the $i_{th}$ label and $FP_{i}$ is the number of false positives for the $i_{th}$ label.
3.6.3 F1-Score
Micro-averaged F1-Score is the harmonic mean of micro-averaged precision and micro-averaged recall, providing a balance between the two. It is particularly useful for evaluating models on imbalanced datasets. The formula for the micro-averaged F1-score is given by eq. (8):
[…]
3.6.4 Jaccard Accuracy
Jaccard Accuracy measures the average similarity between the predicted and ground truth label sets. It is defined as the size of the intersection divided by the size of the union of the predicted and true label sets. The formula for Jaccard accuracy is given by eq. (9):
Jaccard $\text{Jaccard Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{Y}_{i}\cap Y_{i}|}{|\hat{Y}_{i}\cup Y_{i}|}$ (9)
where $N$ is the number of instances, $Y^{i}$ is the predicted label set for instance ii, and $Yi$ is the true label set for instance $i$. The numerator $Y^{i}\cap Yi$ represents the size of the intersection of the predicted and true label sets, and the denominator $Y^{i}\cup Yi$ represents the size of their union.
3.6.5 Hamming Loss
Hamming loss is the fraction of incorrect labels to the total number of labels, providing a measure of how many times an incorrect label is assigned. The formula for Hamming loss is given by eq. (10):
Hamming $\text{Hamming Loss}=\frac{1}{N\cdot C}\sum_{i=1}^{N}\sum_{j=1}^{C}I(\hat{Y}_{ij}\neq Y_{ij})$ (10)
where $N$ is the number of instances, $C$ is the number of classes, $\hat{Y}_{ij}$ is the predicted value for class $j$ of instance $i$, $Y_{ij}$ is the true value for class $j$ of instance $i$, and $I$ is the indicator function that returns $1$ if the predicted and true values are different and 00 otherwise.
These evaluation metrics—precision, recall, F1-score, Jaccard accuracy, and Hamming loss—provide a comprehensive assessment of the model’s performance in multi-label emotion classification for Arabic text, highlighting its strengths and areas for improvement.
4 Results & Discussion
This section presents the experimental results obtained from evaluating the proposed MLEC model for Arabic text. Extensive experiments are conducted to assess the performance of proposed model across several datasets, applying the hybrid loss function to handle class imbalance and label correlations effectively. The results demonstrate the effectiveness of proposed approach, which combines stacked embeddings extraction, hybrid loss and the meta-learner. The results showcase improvements in key evaluation metrics such as precision, recall, F1-score, Jaccard accuracy, and Hamming loss. Through a detailed analysis of these metrics, we provide insights into the strengths and limitations of our model, highlighting its capability to advance the state-of-the-art in Arabic MLEC.
4.1 Experimental Results
Table 3 presents the experimental results evaluated against five metrics: Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. The results compare the baseline method, which uses standard categorical cross-entropy as the loss function, with four variants: Class Weighting (CW), Label Correlation Matrix (LCM), Contrastive Class Weights (CL), and the proposed Hybrid Loss.
Table 3: Evaluation of multi-label emotional classification on SemEval-2018 Task 1-Ec-Ar dataset
Metric Baseline CW LCM CL Hybrid Loss
Precision 0.72 0.75 0.77 0.82 0.82
Recall 0.68 0.72 0.74 0.79 0.81
Micro F1 0.70 0.73 0.75 0.81 0.81
Macro F1 0.70 0.73 0.75 0.81 0.81
Jaccard Accuracy 0.55 0.58 0.61 0.66 0.67
Hamming Loss 0.22 0.20 0.18 0.17 0.15
The results demonstrate that the proposed Hybrid loss based method has good multi-label emotion classification performance. Performance across various metrics indicate that the proposed method is an appropriate choice for the problem at hand, whereas further evaluation may be required to assess the class-wise performance and comparison with the state of the art. It is evident from the results of Table 3 that the proposed Hybrid loss along with the stacked embeddings and meta-learner provides best performance and some of its key features across various metrics are discussed below:
  • Precision: Proposed approach achieves a precision of 0.82, outperforming the baseline (0.72) and showing significant improvement over CW (0.75), LCM (0.77), and matching CL (0.82). This indicates that the proposed approach effectively reduces the number of false positives.
  • Recall: Proposed approach achieves the highest recall of 0.81, which surpasses the baseline (0.68), CW (0.72), LCM (0.74), and CL (0.79). This shows the efficiency of proposed approach in identifying true positives, reducing the number of false negatives.
  • F1-Score: With an F1-score of 0.81, proposed approach shows notable improvement from the baseline (0.70) and other methods (CW: 0.73, LCM: 0.75, CL: 0.81). The high F1-score indicates a balance between precision and recall, demonstrating the model’s robustness.
  • Jaccard Accuracy: Proposed approach achieves the highest JaccardJaccard accuracy of 0.67, outperforming the baseline (0.55), CW (0.58), LCM (0.61), and CL (0.66). This metric underscores the capability of proposed approach to correctly predict all labels for each instance more frequently than other methods.
  • Hamming Loss: Proposed approach has the lowest Hamming loss of 0.15, which is an improvement over the baseline (0.22), CW (0.20), LCM (0.18), and CL (0.17). The lower Hamming loss indicates fewer incorrect label assignments, highlighting the model’s accuracy in predicting multiple labels for each instance.
The findings illustrate that the proposed approach based on Hybrid loss effectively addresses the challenges of class imbalance and label correlation in multi-label classification tasks. By leveraging stacked embeddings, mete-lerarning and hybrid loss based on CW, LCM and CL, the proposed approach provides enhanced performance against various metrics.
4.1.1 Comparison with State-of-the-Art
In this section, we compare the results of our proposed model with existing state-of-the-art models in Arabic MLEC. Table 4 summarizes the performance of these models using precision, recall, micro & macro F1-scores and accuracy. The proposed model demonstrates superior performance compared to the existing state-of-the-art models across multiple metrics. Notably, the proposed model achieves a Micro-Averaged F1-Score of 0.81 and a Jaccard Accuracy of 0.67, which are significantly higher than those of previously reported models.
Table 4: Performance comparison on SemEval-2018 Task 1-Ec-Ar dataset with State-of-the-Art methods
Approach Precision Recall Micro F1 Macro F1 Accuracy
Elfaik et al.[24] - - 0.3204 - 0.5382
Amrita[65] - - 0.379 0.250 0.254
Abdelali et al.[66] - - - 0.468 -
Samy et al.[27] - - 0.495 0.648 0.532
Team-CEN[67] - - 0.516 0.384 0.380
Team-UNCC[68] - - 0.572 0.447 0.446
HEF[26] - - 0.631 0.502 0.512
TW-Star[69] - - 0.597 0.446 0.465
Khalil et al.[23] 0.695 0.551 0.615 0.440 0.498
EMA[70] - - 0.618 0.461 0.489
Mansy et al.[22] 0.634 0.550 0.527 0.701 0.540
DF[26] - - 0.627 0.490 0.505
HEF[26] - - 0.583 0.433 0.448
HEF+DF[26] - - 0.631 0.502 0.512
EmoGraph[71] - - 0.650 0.477 0.523
SpanEmo[72] - - 0.6581 0.5363 0.5394
JSPCL[16] - - 0.6586 0.5370 0.5398
JSCL[16] - - 0.6600 0.5408 0.5414
ICL[16] - - 0.6613 0.5426 0.5417
SLCL[16] - - 0.6637 0.5488 0.5465
SCL[16] - - 0.6671 0.5427 0.5543
Proposed 0.82 0.81 0.81 - 0.67
4.1.2 Class-wise Performance
In this section, we provide a detailed analysis of the performance for each class individually, taking into account the number of samples in each class as shown in Table 5. The dataset exhibits class imbalance, a common challenge in multi-label classification tasks. The issue of class imbalance is significantly reduced by the proposed MLEC due to utilization of hybrid loss, which is designed to mitigate its effects but it still has an impact on class-wise performance as shown in The Table 5.
Table 5: Class-wise performance of proposed approach on SemEval-2018 Task 1-Ec-Ar dataset
Class Name No. of Instances Precision Recall F1-Score Jaccard Accuracy Hamming Loss
Anger 899 0.80 0.78 0.79 0.65 0.12
Anticipation 206 0.74 0.71 0.72 0.60 0.18
Disgust 433 0.76 0.74 0.75 0.62 0.17
Fear 391 0.75 0.73 0.74 0.61 0.18
Joy 605 0.84 0.82 0.83 0.70 0.10
Love 562 0.82 0.80 0.81 0.68 0.11
Optimism 561 0.81 0.79 0.80 0.67 0.12
Pessimism 499 0.79 0.77 0.78 0.64 0.13
Sadness 842 0.77 0.75 0.76 0.66 0.14
Surprise 47 0.70 0.67 0.68 0.50 0.20
Trust 120 0.72 0.70 0.71 0.55 0.18
Neutral 15 0.65 0.62 0.63 0.45 0.22
The results presents the performance metrics for each class, including Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. This detailed analysis evaluates the extent of class imbalance in the dataset and assesses the effectiveness of the use of hybrid loss in the proposed model for addressing these challenges.
4.2 Ablation Study
The ablation study aims to validate the contribution of each component within the proposed model architecture. This study involves evaluating the impact of different loss functions and backbone architectures on the model’s performance. Specifically, we experimented with five variations of the loss function and four different backbone architectures for embedding extraction.
4.2.1 Loss Function Variations
We conducted experiments using the proposed Hybrid loss functions and its variants which includes:
  1. 1.
    Baseline: Standard categorical cross-entropy loss.
  2. 2.
    CW: Incorporates class weights to address class imbalance.
  3. 3.
    LCM: Utilizes a label correlation matrix to account for label dependencies.
  4. 4.
    CL: Employs contrastive learning techniques to improve the discriminative power of the model.
  5. 5.
    Hybrid: The proposed loss function, combining class weights, label correlation, and contrastive learning.
The results of Table 6 demonstrate that MarBERT has the highest performance across all metrics, with a precision of 0.80, recall of 0.78, and an F1-score of 0.79. It also achieved the highest Jaccard accuracy (0.63) and the lowest Hamming loss (0.16). ArabicBERT and AraBERT also showed strong performance, but MarBERT outperformed them in all metrics, indicating its superior ability to capture the nuances of Arabic text.
4.2.2 Backbone Architecture Variations
For embedding extraction, we evaluated three pre-trained language models and each of these models were fine-tuned for the task of MLEC using the Hybrid loss function.
  1. 1.
    ArabicBERT
  2. 2.
    MarBERT
  3. 3.
    AraBERT
The results of Table 6 demonstrate that the Baseline method, using standard categorical cross-entropy loss, performed the worst across all metrics. Introducing CW improved the performance, particularly in terms of precision and recall. Utilizing the LCM further enhanced the model’s performance, highlighting the importance of accounting for label dependencies. CL showed significant improvements, achieving an F1-score of 0.81 and reducing Hamming loss to 0.17. The proposed approach achieved the best overall performance, with an F1-score of 0.81, Jaccard accuracy of 0.67, and the lowest Hamming loss of 0.15. This underscores the effectiveness of combining class weights, label correlations, and contrastive learning. The ablation study confirms that the proposed approach in conjunction with Hybrid loss function provides superior performance for multi-label emotion classification in Arabic text, effectively addressing class imbalance and label correlations.
Table 6: Comparison of Metrics for Different Models
Metric Precision Recall F1-Score Jaccard Accuracy Hamming Loss
Backbone Architecture
ArabicBERT 0.78 0.75 0.76 0.60 0.18
MarBERT 0.80 0.78 0.79 0.63 0.16
AraBERT 0.79 0.76 0.77 0.61 0.17
Loss Function
Baseline 0.72 0.68 0.70 0.55 0.22
CW 0.75 0.72 0.73 0.58 0.20
LCM 0.77 0.74 0.75 0.61 0.18
CL 0.82 0.79 0.81 0.66 0.17
Proposed MLEC 0.82 0.81 0.81 0.67 0.15
4.3 Discussion
The experimental results of Table 3 provide some interesting highlights about the performance of the proposed model which are highlighted below:
  • Precision: Proposed approach achieves a precision of 0.82, outperforming the baseline (0.72) and showing significant improvement over CW (0.75), LCM (0.77), and matching CL (0.82). This indicates that proposed model effectively reduces the number of false positives.
  • Recall: Proposed model achieves the highest recall of 0.81, which surpasses the baseline (0.68), CW (0.72), LCM (0.74), and CL (0.79). This shows efficiency of proposed model in identifying true positives, reducing the number of false negatives.
  • F1-Score: With an F1-score of 0.81, proposed model shows notable improvement from the baseline (0.70) and other methods (CW: 0.73, LCM: 0.75, CL: 0.81). The high F1score indicates a balance between precision and recall, demonstrating the model’s robustness.
  • Jaccard Accuracy: Proposed model achieves the highest Jaccard accuracy of 0.67, outperforming the baseline (0.55), CW (0.58), LCM (0.61), and CL (0.66). This metric underscores capability of propose model to correctly predict all labels for each instance more frequently than other methods.
  • Hamming Loss: Proposed model has the lowest Hamming loss of 0.15, which is an improvement over the baseline (0.22), CW (0.20), LCM (0.18), and CL (0.17). The lower Hamming loss indicates fewer incorrect label assignments, highlighting the model’s accuracy in predicting multiple labels for each instance.
The findings illustrate that proposed model based on hybrid loss effectively addresses the challenges of class imbalance and label correlation in multi-label emotion classification task for Arabic language.
4.3.1 Comparative Performance
Comparison of the proposed approach with state-of-the-art indicate that the proposed approach is superior to existing approaches. The improvements in performance can be attributed to several key factors:
  • Stacked Embeddings: The proposed approach rely on stacked embedding in which the contextualized embeddings extracted from three pre-trained models for Arabic language are used to extract embeddings which are stacked to improve the model’s generalization.
  • Meta-Learning: The stacked embeddings obtained from three language models is provided to a second learner (Bi-LSTM) which performs sequence learning and subsequently predicts the multi-label class vector using a fully connected network.
  • Handling Class Imbalance: The hybrid loss used in the proposed approach incorporates class weighting to address the imbalance present in the dataset. This helps in giving more importance to minority classes, which in turn improves the recall and F1-score for these classes.
  • Leveraging Label Correlations: By integrating a label correlation matrix, hybrid loss effectively captures dependencies between labels, enhancing the overall prediction accuracy. This is particularly beneficial in multi-label classification where labels are often interrelated.
  • Contrastive Learning: The use of contrastive learning further refines the model’s ability to distinguish between different classes by learning from both similar and dissimilar instances. This contributes to higher precision and recall values.
  • Robust Loss Function Design: The enhanced design of proposed hybrid loss function, which integrates class weighting, label correlations, and contrastive learning, provides a comprehensive mechanism to tackle the challenges in multi-label emotion classification.
As a result, the proposed model significantly outperforms existing models mentioned in Table 4, particularly in terms of F1-Score and Jaccard Accuracy. These improvements underscore the effectiveness of our approach in addressing class imbalance and leveraging label correlations, leading to enhanced classification performance in the domain of Arabic emotion classification.
4.3.2 Class-Wise Performance
During analysis of class-wise performance evaluation, it was observed that the model perform with consistently improved performance across various metrics for both majority and minority classes. Some of the key insights are listed below:
  1. 1.
    Majority Classes
    Anger, Joy, Sadness: These classes, with a relatively high number of samples, show strong performance metrics. The high precision, recall, and F1-scores indicate that the model effectively captures these emotions due to the abundant training data.
  2. 2.
    Minority Classes
    Anticipation, Surprise, Trust, Neutral: Despite having fewer samples, the proposed approach significantly enhances the performance for these classes. The precision and recall values, although slightly lower than those of majority classes, demonstrate the model’s improved ability to learn from limited data.
4.3.3 Impact of Hybrid loss
The use of hybrid loss minimize the effects of class imbalance and label correlation which has led to a balanced performance across both majority and minority classes. The reduced Hamming Loss and improved Jaccard Accuracy across most classes highlight the effectiveness of the proposed method.
The class-wise performance analysis underscores the efficacy of the proposed approach in addressing class imbalance issues inherent in multi-label classification tasks. By significantly reducing the performance disparity between majority and minority classes, the proposed approach ensures a more balanced and robust classification model, capable of accurately predicting a wide range of emotions even in the presence of imbalanced data.
5 Conclusion
This study presents a comprehensive approach to improving multi-label emotion classification (MLEC) for the Arabic language by addressing critical challenges such as class imbalance and label correlations. The proposed model integrates stacked embeddings, meta-learning and formulation of a hybrid loss function by integrating class weighting, label correlation matrix, and contrastive learning to effectively tackle these issues. Through rigorous experimentation and performance evaluation, it is evident that the proposed approach significantly enhances the classification performance across various metrics, including Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. The class-wise performance analysis further demonstrates that the use of hybrid loss successfully reduces the disparity between majority and minority classes, ensuring a balanced and robust model capable of accurately predicting a wide range of emotions. Moreover, the ablation study underscores the contribution of each component of the proposed approach, validating its superiority over baseline and other variant loss functions. The detailed comparison with backbone architectures such as ArabicBERT, MarBERT, and AraBERT confirms the effectiveness of the proposed approach in leveraging these models to achieve state-of-the-art performance in emotion classification tasks. In conclusion, the proposed approach represents a significant advancement in the field of multi-label emotion classification for the Arabic language. It not only addresses the prevalent issues of class imbalance and label correlations but also provides a generalizable framework that can be extended to other languages and domains.
Declarations
  • Funding
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use)
  • Data availability
  • Code availability
  • Author contribution